Price elasticity of demand for a commodity is unity and a household demands 50 units of it when its price is rupees 2 per unit. At what price will the household demands 45 units of the commodity?
1
Expert's answer
2019-07-11T11:09:02-0400

P0 =2 P1 = x

Q0 = 50 Q1 = 45

td = (((45-50)*100)/50) / (((x-2)*100)/2) =

= ((-5*100)/50) / ((x-2)*50) = 1

= (-10) / ((x-2)*50) = 1

= 1/(5x-10) = 1

1 = 5x-10

1+10 = 5x

11 = 5x

x = 11/5 = 2.2

Answer:

The household demands 45 items at a price of 2.2 per unit.